In a right triangle, the hypotenuse has length 14 cm. and another side has length 12 cm. What is the length of the altitude that is perpendicular to the hypotenuse?
Answer by checkley75(3666)

14^2=12^2+X^2
X^2=14^2-12^2
X^2=196-144
X^2=52
X=SQRT52
X=7.2111 cm.